# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v1

@@ 1.6. Permutation invariance
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$. The same invariance holds for last-passage percolation if the weights satisfy either $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ with $a_{i}+b_{j}>0$ for all $i,j$, or $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$ for all $i,j$.
[…]
The following is an extension of the recent result [BEM+25a, Theorem 3.1] to multi-paths and to the positive-temperature case. The technique follows the same outline as developed in [BEM+25a], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns. However, the coupling we use in this proof is different because we treat the weights as periodic outside a large box and apply the periodic Pitman transform.
Theorem 1.11.
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$ and $\tau\mathbf{b}=(b_{\tau(j)})_{j\in\mathbb{Z}}$. The same invariance holds for last-passage percolation if the weights satisfy either $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$ or $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$ for all $i,j$.
Theorems 1.10 and 1.11 are proved in Section 5.1. When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COSZ14, OSZ14, BC14]). A similar type of invariance for last-passage percolation was discovered in the work of Dauvergne [Dau22, Theorem 1.5], motivated by a shift-invariance result previously proved in Borodin, Gorin, and Wheeler [BGW22]. Connections between these results are discussed further in [BEM+25a, Section 3].
1.7. Distribution of the jointly invariant measures for a discrete-time Markov chain
The original motivation for introducing the map $D$ in [CGS24] was to understand the jointly invariant measures for the KPZ equation with periodic noise. This was done by first finding jointly invariant measures for the periodic O’Connell-Yor polymer. As a by-product of the analysis, it was shown that the same measures are jointly invariant for a discrete-time Markov chain (see [CGS24, Section 4]). In this section, we study the zero-temperature analogue of that Markov chain. The state space is $\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}\times\mathbb{R}_{\geq 0}^{\mathbb{Z}_{N}}$, which is preserved by the chain by Lemma A.3. If the state of the Markov chain at time $m$ is $(\mathbf{U}_{1}^{(m)},\mathbf{U}_{2}^{(m)})$, and $\mathbf{W}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\geq 0}$ is independent of the chain up to time $m$, then the time $m+1$ state is defined as
$\bigl{(}\mathbf{U}^{(m+1)}_{1},\mathbf{U}^{(m+1)}_{2}\bigr{)}=\bigl{(}\widetilde{D}(\mathbf{W},\mathbf{U}_{1}^{(m)}),\widetilde{D}(\mathbf{W},\mathbf{U}_{2}^{(m)})\bigr{)}.$ (1.14)
Marginally, $\bigl{(}\mathbf{U}_{1}^{(m)}\bigr{)}_{m\geq 0}$ and $\bigl{(}\mathbf{U}_{2}^{(m)}\bigr{)}_{m\geq 0}$ are each Markov chains with the same dynamics. We call the invariant measures for this coupled Markov chain jointly invariant measures because their joint distribution is preserved under the Markov chain where each component evolves using the same driving weights.
For $\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}$, define
$\mathfrak{s}(\mathbf{X})=\sum_{i\in\mathbb{Z}_{N}}X_{i}.$ (1.15)
For $\theta\in\mathbb{R}$, define
$\mathbb{R}^{\mathbb{Z}_{N}}_{\theta}$ $=\{\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid\mathfrak{s}(\mathbf{X})=\theta\},\quad\text{and}$ (1.16)
$\mathbb{R}^{\mathbb{Z}_{N}}_{\theta,\geq 0}$ $=\{\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}_{\theta}\mid X_{i}\geq 0\text{ for }i\in\mathbb{Z}_{N}\}$
We also define $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta}=\mathbb{R}_{\theta}^{\mathbb{Z}_{N}}\cap\mathbb{Z}^{\mathbb{Z}_{N}}$ and $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta,\geq 0}:=\mathbb{R}^{\mathbb{Z}_{N}}_{\theta,\geq 0}\cap\mathbb{Z}^{\mathbb{Z}_{N}}$. We now state the following theorem.
Theorem 1.12.
Let $\theta_{1},\theta_{2}\geq 0$ and $q\in(0,1)$. At each step $m$, let the driving weights $\mathbf{W}$ in the evolution of the Markov chain (1.14) be an i.i.d. sequence of $\text{Geom}(q)$ random variables. Let $\mathbf{X}_{1}$ be uniformly distributed on the set $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{1},\geq 0}$, and let $\mathbf{X}_{2}$ be uniformly distributed on the set $\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{2},\geq 0}$, then the law of $\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}$ is a jointly invariant measure for the Markov chain. Furthermore, we have the distributional equality
$\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\mathbf{X}_{2}\bigr{)}$ (1.17)
The same result holds when $\mathbf{W}$ is an i.i.d. sequence of $\text{Exp}(\alpha)$ random variables for any $\alpha>0$, if we instead take $\mathbf{X}_{1}$ to be uniformly distributed on $\mathbb{R}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$ and $\mathbf{X}_{2}$ to be uniformly distributed on $\mathbb{R}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$.
Theorems 1.12, along with Theorem 1.13 below, is proved in Section 6. In fact, we prove that the same invariance holds when $\mathbf{X}_{1},\mathbf{X}_{2}$ are sequences of i.i.d. geometric random variables. The uniform distribution is obtained by conditioning on $\mathfrak{s}(\mathbf{X}_{1})$ and $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$. These vector sums are preserved by the map $\widetilde{D}$ and hence also the Markov chain (Lemma A.4), so the extremal invariant measures come when we consider $\mathbf{X}_{1},\mathbf{X}_{2}$ to be uniform over all nonnegative vectors with a given sum. As in [CGS24], we can also consider the joint Markov chain with more than two components. In this case, the invariant measures will be built by iterating the map $\widetilde{D}$. Specifically, the invariant initial conditions will be of the form $(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{3})),\ldots$ For brevity, we include only the case of two components to motivate the next theorem.
In this next result, we give an explicit combinatorial description of these jointly invariant measures from Theorem 1.12. We leave it as an open problem to develop a similar combinatorial description for the jointly invariant measures with more than $2$ components. In analogy with [CGS24], when $\mathbf{X}_{1}$ is uniform on $\mathbb{R}_{\theta_{1},\geq 0}^{\mathbb{Z}_{N}}$ and $\mathbf{X}_{2}$ is uniform on $\mathbb{R}_{\theta_{2},\geq 0}^{\mathbb{Z}_{N}}$, the law of $(\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))$ should be a jointly invariant measure for Brownian last-passage percolation in a periodic environment of Brownian motions. It can also be shown (see [CGS24, Appendix F]) that as $N\to\infty$, the invariant measure from Theorem 1.12 converges to the jointly invariant measure for i.i.d. exponential last-passage percolation studied in [FS20], which in turn, converges under diffusive scaling to the $2$-component marginal of the stationary horizon [Bus24]–a coupling of Brownian motions with drift that describes the jointly invariant measures for the KPZ fixed point [BSS24].
Below, we take $\theta_{2}>\theta_{1}$, since we have the distributional equality $\bigl{(}\mathbf{X}_{1},\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr{)}\overset{d}{=}\bigl{(}\widetilde{D}(\mathbf{X}_{2},\mathbf{X}_{1}),\mathbf{X}_{2}\bigr{)}$ in Theorem 1.12. When $\theta_{1}=\theta_{2}$, we have $\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=\mathbf{X}_{1}$ via Corollary 6.4 below, so the joint distribution is trivial in this case.
Below, for $\theta_{1}\leq\theta_{2}$, we define the space
$\mathcal{Z}_{\theta_{1},\theta_{2}}:=\{(\mathbf{u}_{1},\mathbf{u}_{2})\in\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{1},\geq 0}\times\mathbb{Z}^{\mathbb{Z}_{N}}_{\theta_{2},\geq 0}\mid\mathbf{u}_{1}\leq\mathbf{u}_{2}\},$ (1.18)
where $\mathbf{u}_{1}\leq\mathbf{u}_{2}$ means $u_{1,i}\leq u_{2,i}$ for all $i\in\mathbb{Z}_{N}$.
Theorem 1.13.
[…]
$\mathbb{P}(\mathbf{X}_{1}=\mathbf{u}_{1},\mathbf{\widetilde{D}}=\mathbf{u}_{2})=\frac{\alpha(\mathbf{u}_{1},\mathbf{u}_{2})}{\binom{\theta_{1}+N-1}{N-1}\binom{\theta_{2}+N-1}{N-1}}$
where $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ is the following integer, split into two cases:
  1. (i)
    If $u_{1,i}<u_{2,i}$ for all $i\in\mathbb{Z}_{N}$, then $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})=1$.
  2. (ii)
[…]
    $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})=\prod_{n=1}^{s}\sum_{a_{1}=-u_{1,\ell_{m}}}^{0}\sum_{a_{2}=-u_{1,\ell_{m}+1}}^{-a_{1}}\sum_{a_{3}=-u_{1,\ell_{m}+2}}^{-(a_{1}+a_{2})}\cdots\sum_{a_{k_{m+1}-\ell_{m}+N\mathbf{1}_{m=s}}=-u_{1,k_{m+1}-1+N1_{m=s}}}^{-(a_{1}+a_{2}+\cdots+a_{k_{m+1}-\ell_{m}-1+N1_{m=s}})}1.$
    Alternatively, $\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ is the number of $s$-tuples of vectors $(\mathbf{a}_{1},\mathbf{a}_{2},\ldots,\mathbf{a}_{s})$, where, for $m\in\llbracket 1,s\rrbracket$, the vector $\mathbf{a}_{m}$ lies in the set $\mathbb{Z}^{k_{m+1}-\ell_{m}+N1_{m=s}}$ and satisfies the following conditions:
    $a_{m,1}\leq 0$ (1.19)
    $a_{m,1}+a_{m,2}\leq 0$
    $\vdots$
    $a_{m,1}+a_{m,2}+\cdots+a_{m,k_{m+1}-\ell_{m}+N1_{m=s}}\leq 0,\text{ and}$
    $a_{m,p}\geq-u_{1,\ell_{m}+p-1}\text{ for }1\leq p\leq k_{m+1}-\ell_{m}+N1_{m=s}.$
[…]
Figure 5. Comparison of $\mathbf{u}_{1}$ to $\mathbf{u}_{2}$ with $s$ alternating equality/inequality segments. We place an equal sign at index $i$ if $u_{1,i}=u_{2,i}$ and an inequality sign at index $j$ if $u_{1,j}<u_{2,j}$.
Example 1.14.
We demonstrate a sample application of Theorem 1.13 in the case of $N=2$ and $\theta_{1},\theta_{2}\in\mathbb{Z}_{\geq 0}$ where $\theta_{1}<\theta_{2}$. Take some $x_{1},x_{2},y_{1},y_{2}\in\mathbb{Z}_{\geq 0}$ satisfying $x_{1}\leq y_{1},x_{2}\leq y_{2}$ and $x_{1}+x_{2}=\theta_{1},y_{1}+y_{2}=\theta_{2}$.
If $x_{1}=y_{1}$, then $x_{2}<y_{2}$. In the notation of Theorem 1.13, $s=1,k_{1}=1$, and $\ell_{1}=2$, and the only relevant condition is $x_{2}=-u_{1,2}\leq a_{1,1}\leq 0,$ giving $x_{2}+1$ possible values for $a_{1,1}$. We can symmetrically describe the possible values when $x_{1}<y_{1},x_{2}=y_{2}$. Then, we observe that
$\mathbb{P}(\mathbf{X}_{1}=(x_{1},x_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=(y_{1},y_{2}))=\begin{cases}\frac{x_{1}+1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}<y_{1}\text{ and }x_{2}=y_{2}\\ \frac{x_{2}+1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}=y_{1}\text{ and }x_{2}<y_{2}\\ \frac{1}{(\theta_{1}+1)(\theta_{2}+1)}&x_{1}<y_{1}\text{ and }x_{2}<y_{2}\end{cases}$
[…]
Example 1.15.
We similarly demonstrate a computation in the case of $N=3$ and $\theta_{1}<\theta_{2}\in\mathbb{Z}_{\geq 0}$. Take $x_{1},x_{2},x_{3},y_{1},y_{2},y_{3}\in\mathbb{Z}_{\geq 0}$ satisfying $x_{1}\leq y_{1}$,  $x_{2}\leq y_{2},\ x_{3}\leq y_{3}$ and $x_{1}+x_{2}+x_{3}=\theta_{1},\ y_{1}+y_{2}+y_{3}=\theta_{2}$. Write $\mathbf{u}_{1}=(x_{1},x_{2},x_{3}),\mathbf{u}_{2}=(y_{1},y_{2},y_{3})$. Then, we consider the cases in which the only equality is $x_{1}=y_{1}$ and when the only equalities are $x_{1}=y_{1},\ x_{2}=y_{2}$. We can achieve all other sets of equalities and inequalities through cyclic permutations of $\mathbf{u}_{1},\mathbf{u}_{2}$ (Lemma A.6).
When we have that $x_{1}=y_{1},\ x_{2}<y_{2},\ x_{3}<y_{3}$, then we have that $k_{1}=1,\ell_{1}=3$, so the conditions (1.19) are exactly $x_{3}=-u_{1,3}\leq a_{1,1}\leq 0$, giving $x_{3}+1$ distinct possible values. If $x_{1}=y_{1},\ x_{2}=y_{2},\ x_{3}<y_{3}$, then we have that $k_{1}=2,\ell_{1}=3$, so the conditions (1.19) are $x_{3}=-u_{1,3}\leq a_{1,1}\leq 0,\ a_{1,1}+a_{1,2}\leq 0,\ -x_{1}=-u_{1,1}\leq a_{1,2}$. This gives a count of
$\alpha(\mathbf{u}_{1},\mathbf{u}_{2})$ $=\sum_{a_{1}=-x_{3}}^{0}\sum_{a_{2}=-x_{1}}^{-a_{1}}1=\sum_{a_{1}=-x_{3}}^{0}(x_{1}-a_{1}+1)=(x_{3}+1)(x_{1}+1)+\frac{x_{3}(x_{3}+1)}{2}.$
Noting that there are $\binom{\theta_{1}+2}{2}\binom{\theta_{2}+2}{2}$ elements in $\mathbb{Z}_{\theta_{1},+}^{\mathbb{Z}_{N}}\times\mathbb{Z}_{\theta_{2},+}^{\mathbb{Z}_{N}}$, we have the computation
($\binom{\theta_{1}+2}{2}\binom{\theta_{2}+2}{2}\cdot\mathbb{P}$( $\mathbf{X}_{1}=(x_{1},x_{2},x_{3}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})=(y_{1},y_{2},y_{3}))$
$=\begin{cases}(x_{3}+1)(x_{1}+1)+\frac{x_{3}(x_{3}+1)}{2}&x_{1}=y_{1},x_{2}=y_{2},\text{ and }x_{3}<y_{3}\\ \hfil x_{3}+1&x_{1}=y_{1},x_{2}<y_{2},\text{ and }x_{3}<y_{3}\\ \hfil 1&x_{1}<y_{1},x_{2}<y_{2},\text{ and }x_{3}<y_{3}\end{cases}$
@@ 1.8. Organization of the paper
Section 2 is focused on proving the partition function invariance in Theorem 1.4. The braid relations in Theorem 1.6 are proved in Section 3. The details of the zero-temperature limit transitions are handled in Section 4. Section 5 proves inhomogeneous Burke properties for the positive and zero-temperature periodic Pitman transforms and then uses those to prove the permutation invariance in Theorems 1.10 and 1.11. Theorems 1.12 and 1.13 are proved in Section 6. The Appendix contains some prior results from the work [CGS24] and some auxiliary technical results.
@@ 2.1. Matrix encoding
In order to prove our polymer partition function invariance, we follow the matrix encoding method introduced at Noumi and Yamada in [NY04]. We will work with infinite matrices.
@@
First, we justify that this matrix product even makes sense, in that it is finite valued. Because these $H$ matrices (as well as the $E$ matrices) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices. While each row and column has infinitely many non-zero terms, there are only a finite number of non-zero terms in the dot product. Further, the product is upper triangular, so by repeating this argument, we can see that an arbitrary product of $H$ (or $E$) matrices is finite valued.
@@
$=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1}))\qed$
@@ 2.2. Relation to the periodic polymer
$=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
@@
$\sum\limits_{r=i}^{j}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
@@
$Z^{\mathbf{W}}(\ell+1,j\mid k,i)$ $=\sum\limits_{r=-\infty}^{\infty}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
@@
$\mathcal{P}_{k}\mathbf{W}=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2,\ldots})$
@@
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)=Z^{\mathbf{W}}(c,d\mid a,b)$
@@
which we claim is preserved by $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d$. In our base case $n=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$. In the first two sub-cases, $\mathbf{W}_{a}\ldots\mathbf{W}_{c}$ are unaffected by $\mathcal{P}_{k_{1}}$, so our invariance holds trivially.
@@
$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)$
@@
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\sigma}}\ \forall\sigma\in S_{m}}\operatorname{sgn}(\sigma)\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}$
@@
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}=Z^{\mathbf{W}}(V\mid U)$
@@
$\qquad H\Bigl{(}D\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\bigr{)}\Bigr{)}$ (3.4)
@@
$\qquad H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k}))\Bigr{)}$ (3.5)
@@
$=H\Bigl{(}T\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}$
@@
$=e^{Y_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y_{1,i}}+\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N-1}\Biggr{)}$ (3.6)
@@
$(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,N)=e^{Y^{\prime}_{1,N}+Y^{\prime}_{2,N}}$ (3.9)

### v2

@@ 1.6. Zero-temperature limits
The analogues of Theorems 1.7 and 1.8 hold, when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$ and, instead of log-inverse-gamma weights, the weights satisfy either $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$, or $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$, where $a_{i}+b_{j}>0$ (see Definitions 1.13-1.14 below).
The details of the proofs of these theorems are given in Appendix B.
@@ 1.7. Organization of the paper
Section 2 is focused on proving the partition function invariance in Theorem 1.4. The braid relations in Theorem 1.6 are proved in Section 3. Section 4 proves an inhomogeneous Burke property for the periodic Pitman transform and uses it to prove Theorems 1.7 and 1.8. Appendix A contains some prior results from the work [CGS24] and some auxiliary technical results. The necessary details for the zero-temperature analogues in Theorems 1.10-1.12 are proved in Appendix B.
@@ 2.1. Matrix encoding
In order to prove our polymer partition function invariance, we follow the matrix encoding method introduced at Noumi and Yamada in [NY04]. The key difference is that we will work with infinite matrices.
@@
First, we justify that this matrix product has finite entries. Because these $H$ matrices (as well as the $E$ matrices) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices. While each row and column has infinitely many non-zero terms, there are only a finite number of non-zero terms in the dot product. Further, the product is upper triangular, so by repeating this argument, we can see that an arbitrary product of $H$ (or $E$) matrices is finite valued.
@@
$=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).\qed$
@@ 2.2. Relation to the periodic polymer
$=\sum\limits_{r=i}^{j}\Biggl{(}Z^{\mathbf{W}}(\ell,r\mid k,i)\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
@@
$\sum\limits_{r=i}^{j}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}.$
@@
$\quad\;Z^{\mathbf{W}}(\ell+1,j\mid k,i)$
$=\sum\limits_{r=-\infty}^{\infty}\Biggl{(}\Bigl{(}H(\mathbf{W}_{k})H(\mathbf{W}_{k+1})\ldots H(\mathbf{W}_{\ell})\Bigr{)}_{(i,r)}\Bigl{(}H(\mathbf{W}_{\ell+1})\Bigr{)}_{(r,j)}\Biggr{)}$
@@
$\mathcal{P}_{k}\mathbf{W}=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2,\ldots}).$
@@
$Z^{\mathcal{P}_{k_{m}}\cdots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b)=Z^{\mathbf{W}}(c,d\mid a,b).$
@@
which we claim is preserved by the action $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ for any $b,d\in\mathbb{Z}$. In our base case $n=1$, we split into three sub-cases: $k_{1}<a-1$, $k_{1}>c$, and $a\leq k_{1}<c$. In the first two sub-cases, $\mathbf{W}_{a}\ldots\mathbf{W}_{c}$ are unaffected by $\mathcal{P}_{k_{1}}$, so our invariance holds trivially.
@@
$=Z^{\mathcal{P}_{k_{m}}\mathcal{P}_{k_{m-1}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(c,d\mid a,b).$
@@
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V_{\sigma}}\ \forall\sigma\in S_{m}}\operatorname{sgn}(\sigma)\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}.$
@@
$\det M=\sum\limits_{P=(P_{1}\ldots P_{m})\in\Pi_{U,V}}\prod\limits_{(i,j)\in P}e^{W_{(i,j)}}=Z^{\mathbf{W}}(V\mid U).$
@@
$\qquad H\Bigl{(}D\bigl{(}D(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),D\bigl{(}T(\mathbf{W}_{k+2},\mathbf{W}_{k+1}),\mathbf{W}_{k}\bigr{)}\bigr{)}\Bigr{)}.$ (3.4)
@@
$\qquad H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D(\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k}))\Bigr{)}.$ (3.5)
@@
$=H\Bigl{(}T\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}H\Bigl{(}D\bigl{(}T\bigl{(}\mathbf{W}_{k+2},D(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)},T(\mathbf{W}_{k+1},\mathbf{W}_{k})\bigr{)}\Bigr{)}.$
@@
$=e^{Y_{2,N}}\Biggl{(}\prod_{i=1}^{N}e^{Y_{1,i}}+\left(H(\mathbf{Y}_{1})H(\mathbf{Y}_{2})\right)_{1,N-1}\Biggr{)}.$ (3.6)
@@
$(H(\mathbf{Y}^{\prime}_{1})H(\mathbf{Y}_{2}^{\prime}))_{(N,N)}=Z^{\mathbf{Y}^{\prime}}(2,N\mid 1,N)=e^{Y^{\prime}_{1,N}+Y^{\prime}_{2,N}}.$ (3.9)